The shorter leg of a right triangle is 7 yards. If the hypotenuse is 25 yards, how long is the other leg?
Answer by stanbon(75887)

The shorter leg of a right triangle is 7 yards. If the hypotenuse is 25 yards, how long is the other leg?
----------
Let the longer leg be "x":
x^2 + 7^2 = 25^2
x^2 + 49 = 625
--
x^2 = 576
---
x = 24 yds
=================
Cheers,
Stan H.
===============